OpenStudy (anonymous):

Find the arc length of: y=(x^3/3)+(1/4x) 1

OpenStudy (amistre64):

arc length has to do with ds right?

OpenStudy (amistre64):

\[ds = \sqrt{(x')^2+(y')^2}\]

OpenStudy (anonymous):

Arc length \[\int\limits_{a}^{b} \sqrt{1+[f \prime(x)]^2}\]

OpenStudy (amistre64):

since x' with regard to x = 1 \[ds = \sqrt{1+(y')^2}\]

OpenStudy (amistre64):

y= (x^3/3) + (1/4x) ; is the 4x all a denom?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

x^3/3 derived is: 3/3 x^2 = x^2 1/4 x^(-1) derived is: -1/4 x^(-2)

OpenStudy (amistre64):

\[\left(x^2-\frac{1}{4}x^{-2}\right)^2=x^4-\frac{1}{2}+\frac{1}{16}x^{-4}\]

OpenStudy (amistre64):

add 1 to that and we have: \[\sqrt{x^4+\frac{1}{16x^4}+\frac{1}{2}}\] that doesnt look very nice at all does it

OpenStudy (anonymous):

Then complete the square?

OpenStudy (amistre64):

i was thinking more like opening up a half gallon of happy tracks icecream, balling up in the corner, and crying :)

OpenStudy (anonymous):

There has to be some way to break this down.

OpenStudy (amistre64):

maybe swithing it to polars ?

OpenStudy (anonymous):

I think that's too much work.

OpenStudy (anonymous):

Wouldn't be be easier if we left it as 1+(x^2+(1/16x^2)

OpenStudy (amistre64):

dunno, but i dount it what if we add them first?

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (amistre64):

add the fractions to get a quotient \[\frac{x^3}{3}+\frac{1}{4x}=\frac{4x^4+3}{12x}\]

OpenStudy (anonymous):

Yeah, I forgot.

OpenStudy (amistre64):

\[y'=\frac{12x(16x^3)-12(4x^4+3)}{12*12x^2}\] \[y'=\frac{12x^4+3}{12x^2}\] \[y'=\frac{4x^4+1}{4x^2}\]

OpenStudy (amistre64):

-1, not +1 forgot to carry the negative all the way thru

OpenStudy (anonymous):

that does not look like it's easier

OpenStudy (amistre64):

square that for\[(y')^2=\frac{16x^8-8x^4+1}{(4x^2)^2}\] \[1+(y')^2=\frac{16x^8-8x^4+1+16x^4}{(4x^2)^2}\] \[1+(y')^2=\frac{16x^8+8x^4+1}{(4x^2)^2}\] \[\sqrt{1+(y')^2}=\frac{\sqrt{16x^8+8x^4+1}}{4x^2}\] the top is a quadrquadratic in disguise

OpenStudy (anonymous):

Oh, I see that's a lot easier

OpenStudy (amistre64):

say u = x^2 \[\frac{\sqrt{16(u^4+\frac{1}{2}u^2)+1}}{4u}\] \[\frac{\sqrt{16(u^4+\frac{1}{2}u^2+\frac{1}{16})}}{4u}\] \[\frac{\sqrt{16(u^2+\frac14)^2}}{4u}\] \[\frac{u^2+\frac14}{u}\]

OpenStudy (amistre64):

\[\int_{a}^{b}u+\frac{1}{4u}du\]

OpenStudy (amistre64):

since x = 1 to 2 and u = x^2 then u = 1^2 to 2^2 a = 1, b=4

OpenStudy (amistre64):

i think i missed a math someplace along the way :/

OpenStudy (anonymous):

I'll check

OpenStudy (amistre64):

u = x^2 du = 2x dx dx = du/2sqrt(u) so maybe:\[\int_{1}^{4}\frac{u}{2\sqrt{u}}+\frac{1}{8u\sqrt{u}}du\]

OpenStudy (amistre64):

forgot to change my dx to a proper du part :)

OpenStudy (anonymous):

Can't the first one be sqrtu/2?

OpenStudy (amistre64):

yes, and that ones chks out better :) i get 2.4583 for each route

OpenStudy (anonymous):

yeah, that's right

OpenStudy (amistre64):

yay!! :)